# Local EGOP for Continuous Index Learning

### v2

@@ B.3.7 Noisy Manifold
We start off by iterating Corollary 1 as described in the text.
@@
Suppose that for all $v\in\operatorname{Null}$, $\eta\in\operatorname{Null}^{\perp}$, $\eta^{T}\Sigma v=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2})$. Then, for $\Sigma_{\operatorname{Null}^{\perp}}:=\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}^{\perp}}$, $\Sigma_{\operatorname{Null}}:=\pi_{\operatorname{Null}}\Sigma\pi_{\operatorname{Null}}$ there exists $H_{\Sigma_{\operatorname{Null}}},T_{\Sigma_{\operatorname{Null}}}$ such that
@@
$\tilde{R}_{C,i}:=\bar{R}_{C,i}-\tfrac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T},\quad\bar{C}_{i}:=\beta C_{i}^{\prime}+(1-\beta)C_{i}^{\prime},\quad A_{i}=\bar{Q}_{\operatorname{Null}^{\perp}}\bar{C}_{i}\bar{Q}_{\operatorname{Null}^{\perp}}+\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}},$
$W_{i}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}-\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}}\bar{[}Q_{\operatorname{Null}^{\perp}}\bar{C}_{i}\bar{Q}_{\operatorname{Null}^{\perp}}]^{-1}\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}.$
$C_{i+1}=t_{i+1}\begin{bmatrix}A_{i}^{-1}+A_{i}^{-1}[\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}]W_{i}^{-1}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}}]A_{i}^{-1}&-A_{i}^{-1}\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}W_{i}^{-1}\\ -W_{i}^{-1}Q_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}^{\perp}}A_{i}^{-1}&W_{i}^{-1}\end{bmatrix}$
@@
Recall that $C_{i+1}:=t_{i+1}(\bar{C}_{i}+\tilde{R}_{C,i})^{-1}$, hence this is the block Schur complement in the specified basis pivoted on the $\bar{Q}_{\operatorname{Null}^{\perp}}\times\bar{Q}_{\operatorname{Null}^{\perp}}$ block. ∎
@@
The matrix $\alpha^{2}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}]^{-1}=\Theta(1)$, $i=0,1$.
@@
Suppose that $X_{0}=\alpha I$ and $t_{1}=\alpha^{2}$. In the basis $(g,\overline{Q}_{\operatorname{Null}^{\perp}},\overline{Q}_{\operatorname{Null}}),$ we have
@@
$\Sigma_{i+1}=t_{i+1}[\beta\mathcal{L}(\mu_{i}(\zeta))+(1-\beta)\mathcal{L}(\mu_{i-1}(\zeta)]^{-1}$
$\sqrt{t_{i}}=\alpha r^{i}$, $0<r<1,\ \beta>0$. Then, for $\alpha,\zeta$ sufficiently small, $g^{T}\Sigma_{i}g=\Theta(t_{i})$, for $v\in\operatorname{Null}$, $v^{T}\Sigma_{i}v=\Theta(1)$, and for $\eta\in\operatorname{Null}^{\perp}\cap g^{\perp}$, $\eta^{T}\Sigma\eta=\theta(\sqrt{t})$.
@@
$\alpha_{i}=\Theta(t_{i}),\quad b_{i}=O(\sqrt{t_{i}}),\quad C_{i}=\Theta(\sqrt{t_{i}}),\quad\bar{R}_{C,i}=O(t_{i}),$
$\bar{R}_{b,i}=O(\sqrt{t_{i}}),\quad\bar{R}_{\alpha,i}=O(\sqrt{t_{i}}),\quad\|\bar{Q}_{\operatorname{Null}^{\perp}}\Sigma\bar{Q}_{\operatorname{Null}}\|=O(\sqrt{t_{i}}).$
The weak $\operatorname{Null}$, $\operatorname{Null}^{\perp}$ corelation allows us to apply Lemma 26 to reduce the recurrence to the marginal form
@@
$A_{i}=H_{\bar{Q}_{\operatorname{Null}^{\perp}}}[\beta C_{i}+(1-\beta)C_{i-1}]H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(t_{i}),\quad W_{i}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}+O(t_{i}^{3/2}).$
@@
$\bar{Q}_{\operatorname{Null}^{\perp}}C_{i+1}\bar{Q}_{\operatorname{Null}^{\perp}}=t_{i+1}\{H_{\bar{Q}_{\operatorname{Null}^{\perp}}}[\beta(\bar{Q}_{\operatorname{Null}^{\perp}}C_{i}\bar{Q}_{\operatorname{Null}^{\perp}})+(1-\beta)(\bar{Q}_{\operatorname{Null}^{\perp}}C_{i-1}\bar{Q}_{\operatorname{Null}^{\perp}})]H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(t_{i})\}^{-1}.$
by the reductions from Propostions 2 and 3, we can reduce this to an autonomous recurrence in $Q_{\operatorname{Null}^{\perp}}C_{i}\bar{Q}_{\operatorname{Null}^{\perp}}$ that converges at the desired rate by Lemma 23. See the the proof of Theorem 1 for more details.
@@
$\bar{Q}_{\operatorname{Null}}C_{i+1}\bar{Q}_{\operatorname{Null}}=t_{i+1}W_{i}^{-1}=t_{i+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})^{-1}+O(\sqrt{t_{i}}).$
As $\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}}=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2})=O(t_{i})$, it follows that $t_{i+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})^{-1}=\Omega(1)$, and thus it is of constant order by our covariance cap. Thus this rate is an immediate consequence of the second order anisotropy, and the orthogonal component and other remainders only contribute negligibly, justifying the uniform rate. The remaining terms in the Lemma 27 block decomposition are
$\bar{Q}_{\operatorname{Null}^{\perp}}\Sigma_{i+1}\bar{Q}_{\operatorname{Null}}=\bar{Q}_{\operatorname{Null}^{\perp}}C_{i+1}\bar{Q}_{\operatorname{Null}}=-t_{i+1}A_{i}^{-1}(\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})W_{i}^{-1},\quad b_{i+1}=u^{T}\Sigma_{i+1}\bar{Q}=C_{i+1}(b_{i}+R_{b,i}),$
for $u=g/\|g\|$ the normalized gradient vector. In the first term, the product $(\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,i}\bar{Q}_{\operatorname{Null}})W_{i}^{-1}=\Theta(1)$, hence the rate is dominated by the second order anisotropy $t_{i+1}A_{i}^{-1}=\Theta(\sqrt{t_{i}})$. For the final vector, $\|b_{i}+R_{b,i}\|=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|)$, hence we can combine this with our previous analysis to yield
$\bar{Q}_{\operatorname{Null}}C_{i+1}(b_{i}+R_{b,i})=\Theta(1)O(\sqrt{t_{i}}))=O(\sqrt{t_{i+1}}),\quad\bar{Q}_{\operatorname{Null}^{\perp}}C_{i+1}(b_{i}+R_{b,i})=\Theta(t_{i+1})O(\sqrt{t_{i}})=O(t_{i+1}).$
Note that the scalar inflation to go from $t_{i}$ to $t_{i+1}$ does not aggregate across iterations as the initial bound was driven by the second order anisotropy term, to which the remainders only contribute negligibly.
@@
As $\zeta\to 0$, $H_{\zeta}\to H$, as the distribution concentrates about $x^{*}$, and in particular, for $\zeta$ small enough, the signature of $H_{\zeta}$ is the same as that of $H$. This follows from the maximal rank assumption on $H$. By Propositions 2 and 3, upon transforming $C_{i,\zeta}\mapsto\frac{W_{\zeta}^{T}C_{i,\zeta}W_{\zeta}}{\sqrt{c(r,\beta)t_{i}}}=:X_{i,\zeta}$, $W_{\zeta}:=U_{\zeta}|\Lambda_{\zeta}|^{1/2}$, for the diagonalization $H_{\zeta}=U_{\zeta}\Lambda_{\zeta}U_{\zeta}^{T}$, these recurrences are reduced to the same canonical form, with seeds depending smoothly on $\zeta$. Applying Lemma 21, it follows that $X_{i,\zeta}=X_{i,0}+O(\zeta+\alpha)$.
Inverting the transformation, this yields $C_{i,\zeta}=C_{i,0}+O(\sqrt{t_{i}}(\delta+\alpha)).$ Returning to the proof of Lemma 28,
$\Sigma_{i,\operatorname{Null},\zeta}=t_{i+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,i,\zeta}(\Sigma_{i,\operatorname{Null}^{\perp},\zeta})\bar{Q}_{\operatorname{Null}})^{-1}+O(\sqrt{t_{i}})=\Sigma_{i,\operatorname{Null}^{\perp},0}+O(\sqrt{t_{i}}(\zeta+\alpha)),$

### v3

@@ B.3.7 Noisy Manifold
We start off by iterating Corollary 1 as described in the text. Let $\Sigma_{\operatorname{Null}^{\perp}}:=\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}^{\perp}}$, $\Sigma_{\operatorname{Null}}:=\pi_{\operatorname{Null}}\Sigma\pi_{\operatorname{Null}}$.
@@
Suppose that for all $v\in\operatorname{Null}$, $\eta\in\operatorname{Null}^{\perp}$, $\eta^{T}\Sigma v=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|)$, and $\Sigma_{\operatorname{Null}}=\Theta(1)$. There exists $H_{\Sigma_{\operatorname{Null}}},T_{\Sigma_{\operatorname{Null}}}$ such that
@@
$\tilde{R}_{C,t}:=\bar{R}_{C,t}-\tfrac{1}{\eta^{2}+\alpha_{t}+\bar{R}_{\alpha,t}}\bar{b}_{t}\bar{b}_{t}^{T},\quad\bar{C}_{t}:=\beta C_{t}^{\prime}+(1-\beta)C_{t}^{\prime},\quad A_{t}=\bar{Q}_{\operatorname{Null}^{\perp}}\bar{C}_{t}\bar{Q}_{\operatorname{Null}^{\perp}}+\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}^{\perp}},$
$W_{t}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}-\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}^{\perp}}\bar{[}Q_{\operatorname{Null}^{\perp}}\bar{C}_{t}\bar{Q}_{\operatorname{Null}^{\perp}}]^{-1}\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}.$
$C_{t+1}=h_{t+1}\begin{bmatrix}A_{t}^{-1}+A_{t}^{-1}[\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}]W_{t}^{-1}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}^{\perp}}]A_{t}^{-1}&-A_{t}^{-1}\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}W_{t}^{-1}\\ -W_{t}^{-1}Q_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}^{\perp}}A_{t}^{-1}&W_{t}^{-1}\end{bmatrix}$
@@
Recall that $C_{t+1}:=t_{t+1}(\bar{C}_{t}+\tilde{R}_{C,t})^{-1}$, hence this is the block Schur complement in the specified basis pivoted on the $\bar{Q}_{\operatorname{Null}^{\perp}}\times\bar{Q}_{\operatorname{Null}^{\perp}}$ block. ∎
@@
The matrix $\alpha^{2}[\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}]^{-1}=\Theta(1)$, $t=0,1$.
@@
Suppose that $X_{0}=\alpha I$ and $h_{1}=\alpha^{2}$. In the basis $(g,\overline{Q}_{\operatorname{Null}^{\perp}},\overline{Q}_{\operatorname{Null}}),$ we have
@@
$\Sigma_{t+1}=t_{t+1}[\beta\mathcal{L}(\mu_{t}(\zeta))+(1-\beta)\mathcal{L}(\mu_{t-1}(\zeta)]^{-1}$
$\sqrt{t_{t}}=\alpha r^{t}$, $0<r<1,\ \beta>0$. Then, for $\alpha,\zeta$ sufficiently small, $g^{T}\Sigma_{t}g=\Theta(h_{t})$, for $v\in\operatorname{Null}$, $v^{T}\Sigma_{t}v=\Theta(1)$, and for $\eta\in\operatorname{Null}^{\perp}\cap g^{\perp}$, $\eta^{T}\Sigma\eta=\theta(\sqrt{h_{t}})$.
@@
$\alpha_{t}=\Theta(h_{t}),\quad b_{t}=O(\sqrt{h_{t}}),\quad C_{t}=\Theta(\sqrt{h_{t}}),\quad\bar{R}_{C,t}=O(h_{t}),$
$\bar{R}_{b,t}=O(\sqrt{h_{t}}),\quad\bar{R}_{\alpha,t}=O(\sqrt{h_{t}}),\quad\|\bar{Q}_{\operatorname{Null}^{\perp}}\Sigma_{t}\bar{Q}_{\operatorname{Null}}\|=O(\sqrt{h_{t}}).$
The weak $\operatorname{Null}$, $\operatorname{Null}^{\perp}$ correlation allows us to apply Lemma 26 to reduce the recurrence to the marginal form
@@
$A_{t}=H_{\bar{Q}_{\operatorname{Null}^{\perp}}}[\beta C_{t}+(1-\beta)C_{t-1}]H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(h_{t}),\quad W_{t}=\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}+O(h_{t}^{3/2}).$
@@
$\bar{Q}_{\operatorname{Null}^{\perp}}C_{t+1}\bar{Q}_{\operatorname{Null}^{\perp}}=t_{t+1}\{H_{\bar{Q}_{\operatorname{Null}^{\perp}}}[\beta(\bar{Q}_{\operatorname{Null}^{\perp}}C_{t}\bar{Q}_{\operatorname{Null}^{\perp}})+(1-\beta)(\bar{Q}_{\operatorname{Null}^{\perp}}C_{t-1}\bar{Q}_{\operatorname{Null}^{\perp}})]H_{\bar{Q}_{\operatorname{Null}^{\perp}}}+O(h_{t})\}^{-1}.$
by the reductions from Propostions 2 and 3, we can reduce this to an autonomous recurrence in $Q_{\operatorname{Null}^{\perp}}C_{t}\bar{Q}_{\operatorname{Null}^{\perp}}$ that converges at the desired rate by Lemma 23. See the the proof of Theorem 1 for more details.
@@
$\bar{Q}_{\operatorname{Null}}C_{t+1}\bar{Q}_{\operatorname{Null}}=h_{t+1}W_{t}^{-1}=h_{t+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}})^{-1}+O(\sqrt{h_{t}}).$
As $\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}}=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|^{2})=O(h_{t})$, it follows that $h_{t+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}})^{-1}=\Omega(1)$, and thus it is of constant order by our covariance cap. Thus this rate is an immediate consequence of the second order anisotropy, and the orthogonal component and other remainders only contribute negligibly, justifying the uniform rate. The remaining terms in the Lemma 27 block decomposition are
$\bar{Q}_{\operatorname{Null}^{\perp}}\Sigma_{t+1}\bar{Q}_{\operatorname{Null}}=\bar{Q}_{\operatorname{Null}^{\perp}}C_{t+1}\bar{Q}_{\operatorname{Null}}=-h_{t+1}A_{t}^{-1}(\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}})W_{t}^{-1},\quad b_{t+1}=u^{T}\Sigma_{t+1}\bar{Q}=C_{t+1}(b_{t}+R_{b,t}),$
for $u=g/\|g\|$ the normalized gradient vector. In the first term, the product $(\bar{Q}_{\operatorname{Null}^{\perp}}\tilde{R}_{C,t}\bar{Q}_{\operatorname{Null}})W_{t}^{-1}=\Theta(1)$, hence the rate is dominated by the second order anisotropy $h_{t+1}A_{t}^{-1}=\Theta(\sqrt{h_{t}})$. For the final vector, $\|b_{t}+R_{b,t}\|=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|)$, hence we can combine this with our previous analysis to yield
$\bar{Q}_{\operatorname{Null}}C_{t+1}(b_{t}+R_{b,t})=\Theta(1)O(\sqrt{h_{t}}))=O(\sqrt{h_{t+1}}),\quad\bar{Q}_{\operatorname{Null}^{\perp}}C_{t+1}(b_{t}+R_{b,t})=\Theta(h_{t+1})O(\sqrt{h_{t}})=O(h_{t+1}).$
Note that the scalar inflation to go from $h_{t}$ to $h_{t+1}$ does not aggregate across iterations as the initial bound was driven by the second order anisotropy term, to which the remainders only contribute negligibly. ∎
@@
As $\zeta\to 0$, $H_{\zeta}\to H$, as the distribution concentrates about $x^{*}$, and in particular, for $\zeta$ small enough, the signature of $H_{\zeta}$ is the same as that of $H$. This follows from the maximal rank assumption on $H$. By Propositions 2 and 3, upon transforming $C_{t,\zeta}\mapsto\frac{W_{\zeta}^{T}C_{t,\zeta}W_{\zeta}}{\sqrt{c(r,\beta)h_{t}}}=:X_{t,\zeta}$, $W_{\zeta}:=U_{\zeta}|\Lambda_{\zeta}|^{1/2}$, for the diagonalization $H_{\zeta}=U_{\zeta}\Lambda_{\zeta}U_{\zeta}^{T}$, these recurrences are reduced to the same canonical form, with seeds depending smoothly on $\zeta$. Applying Lemma 21, it follows that $X_{t,\zeta}=X_{t,0}+O(\zeta+\alpha)$.
Inverting the transformation, this yields $C_{t,\zeta}=C_{t,0}+O(\sqrt{h_{t}}(\delta+\alpha)).$ Returning to the proof of Lemma 28,
$\Sigma_{t,\operatorname{Null},\zeta}=h_{t+1}(\bar{Q}_{\operatorname{Null}}\tilde{R}_{C,t,\zeta}(\Sigma_{t,\operatorname{Null}^{\perp},\zeta})\bar{Q}_{\operatorname{Null}})^{-1}+O(\sqrt{h_{t}})=\Sigma_{t,\operatorname{Null}^{\perp},0}+O(\sqrt{h_{t}}(\zeta+\alpha)),$